The waiting times on hold for a call to a customer service department are observed to have the following probability distribution Number of calls, x Waiting time, minutes, f(x) 3 10 15 10 6 2 4 10 The expected value of the waiting time for a random call is most nearly: 7.3 minutes 9.2 minutes 11.5 minutes 15.0 minutes
Answer #1

Solution
We are given that: The waiting times on hold for a call to a customer service department are observed to have the following probability distribution:

Number of Calls x Waiting time, minutes f(x)
0 0
2 3
4 10
6 15
8 10
10 6

We have to find expected value of of the waiting time for a random call.

Σ f(r) r E(f(X)) =

Number of Calls x Waiting time, minutes f(x) f(x) * x
0 0 0
2 3 6
4 10 40
6 15 90
8 10 80
10 6 60
\sum x=30 \sum f(x) \times x=276

Thus

E(f(X))=\frac{\sum f(x) \times x}{\sum x}

E(f(X))=\frac{276}{30}

E(f(X))=9.2

Thus correct option is second option = 9.2 minutes.